# Testing the Star-disk Collision Model for Quasi-periodic Eruptions

### v1

@@ 3 Constraints for the Model
$R=\left[GM_{\rm{h}}\left(\frac{P}{2\pi}\right)^{2}\right]^{\frac{1}{3}}=110P_{\rm{1}}^{\frac{2}{3}}M_{\rm{6}}^{-\frac{2}{3}}\ R_{\rm{g}},$ (1)
where $R_{\rm{g}}=\frac{GM_{\rm{h}}}{c^{2}}$ is the gravitational radius of the BH and $c$ is the speed of light in vacuum. In order to simplify the derivation, we define the following notation: $r_{*}=R_{*}/\rm{R}_{\rm{\odot}}$, $M_{\rm{6}}=M_{\rm{h}}/10^{6}\ \rm{M}_{\rm{\odot}}$, $P_{\rm{1}}=P_{\rm{orb}}/10\ \rm{hr}$ and $\alpha_{-1}=\alpha/0.1$, where $\rm{R}_{\rm{\odot}}$ and $\rm{M}_{\rm{\odot}}$ are the solar radius and mass, $\alpha$ is the dimensionless disk viscosity parameter and $R_{*}$ is the radius of the star.
3.1 The Disk
@@
For the inner region of the disk where it is radiation-pressure and radiative-cooling dominated, the scale height of the disk is given by (also see Frank et al., 2002; Linial & Metzger, 2023),
@@
where $v_{k}$ is the Kepler velocity at radius $R$, $\kappa$ is the opacity, $\dot{M}$ is the mass accretion rate at the collision point. Hereafter we consider that the opacity is dominated by Thompson scattering, thus $\kappa=0.34\ \rm{cm}^{2}\rm{g}^{-1}$. The last step of the above was used $R\gg 2R_{\rm{g}}$. Thus, the viscosity can be expressed as $\nu\approx\alpha\sqrt{GMR}(\frac{h}{R})^{2}$. Furthermore, we use the lowercase letter $\dot{m}\equiv\dot{M}/\left(L_{\rm{Edd}}/c^{2}\right)$ to denote the dimensionless accretion rate, where $L_{\rm{Edd}}$ is the Eddington luminosity. Under this definition, the luminosity created by accretion is $L=\eta\dot{m}L_{\rm{Edd}}$, where $\eta$ is the radiation efficiency.
Based on the $\alpha$-disk model (Shakura & Sunyaev, 1973), the surface density is given as
@@
where $\nu\approx\alpha c_{\rm{s}}h$ is the viscosity of the disk.
@@
$M_{\rm{sh}}$ $=2\pi R_{*}^{2}\left(\sqrt{2}\Sigma_{\rm{d}}\right)$
$=4.3\times 10^{-7}r_{*}^{2}\frac{P_{\rm{1}}}{\alpha_{-1}\dot{m}M_{6}}\ \rm{M}_{\rm{\odot}}.$ (4)
The factor of $\sqrt{2}$ comes from the consideration that the velocity of the disk material is perpendicular to the stellar velocity and both are of the order of to the Kepler velocity $v_{\rm{k}}=\sqrt{\frac{GM_{\rm{h}}}{R}}$, so the effective length of the stellar trajectory is $2\sqrt{2}h$.
3.2 Hydrodynamics in Collision
@@ 3.3 Duration of Emission
$\Sigma=\frac{M_{\rm{diff}}}{4\pi R_{\rm{diff}}^{2}}$ (7)
is the surface density of the fireball, $R_{\rm{diff}}=v_{\rm{ej}}t_{\rm{e}}$ is the radius of the fireball at its peak and $t_{\rm{e}}$ is the duration of the eruption. Therefore, one has (also see Linial & Metzger, 2023)
$t_{\rm{e}}\approx\left(\frac{\kappa M_{\rm{sh}}}{4\pi cv_{\rm{ej}}}\right)^{\frac{1}{2}}\approx 0.18\frac{r_{*}}{\alpha_{-1}^{\frac{1}{2}}\dot{m}^{\frac{1}{2}}}\left(\frac{P_{\rm{1}}}{M_{\rm{6}}}\right)^{\frac{2}{3}}\ \rm{hr}.$ (8)
3.4 Luminosity
Next, we consider the energy evolution of the shocked matter. The following Eq. (5), the total internal energy of the immediately post-shock matter, which is dominated by radiation, is $E_{\rm{i,0}}=M_{\rm{sh}}v_{\rm{k}}^{2}$. Note that it is on the same order of magnitude as the total kinetic energy of the shocked matter $E_{\rm{k}}=\frac{1}{2}M_{\rm{sh}}v_{\rm{k}}^{2}$ (Linial & Metzger, 2023).
@@ 3.4 Luminosity
$E_{i}\left(R_{\rm{diff}}\right)$ $=E_{\rm{i,0}}\times\left(\frac{2\pi R_{*}^{2}h/7}{\frac{4\pi}{3}R_{\rm{diff}}^{3}}\right)^{\frac{1}{3}}$
$=3.5\times 10^{44}r_{*}^{\frac{5}{3}}\dot{m}^{-\frac{1}{6}}\alpha_{\rm{-1}}^{-\frac{1}{2}}M_{\rm{6}}^{\frac{1}{3}}\ \rm{erg/s}.$ (9)
@@
$L_{\rm{p}}=\frac{E_{i}\left(R_{\rm{diff}}\right)}{t_{\rm{e}}}=5.5\times 10^{41}M_{\rm{6}}\left(\frac{r_{*}}{P_{\rm{1}}}\right)^{\frac{2}{3}}\dot{m}^{\frac{1}{3}}\ \rm{erg/s}.$ (10)
@@
3.5 The Quiescent-level Accretion
@@
3.6 Infer $r_{*}$ from Observables
Eqs. (8), (10) and (11) give the three observables $(t_{\rm{PQE}}$, $L_{\rm{p}}$ and $L_{\rm{Q}})$ from the star-disk collision model. Here, we attempt to investigate what $r_{*}$ is allowed or required by the model to satisfy these observables.
Notice that $\alpha$ and $\eta$ are the two fudge “factors” that describe the working mechanism of the accretion disk, and $\dot{m}$ is one of the disk properties; all others are observable, or measurable in principle. One can get rid of $\dot{m}$ and express $r_{*}$ in terms of observables in two ways. First, one can infer $\dot{m}$ from Eq. (10), and plug it into Eq. (8) to get
$r_{*}=3.7L_{\rm{p,42}}^{\frac{3}{4}}P_{\rm{1}}^{\frac{1}{6}}M_{\rm{6}}^{-\frac{5}{12}}\left(\frac{t_{\rm{e}}}{1\ \rm{hr}}\right)^{\frac{1}{2}}{\alpha_{-1}}^{\frac{1}{4}}.$ (12)
@@ 3.6 Infer $r_{*}$ from Observables
$r_{*}=28\ L_{\rm{p,42}}^{\frac{3}{2}}L_{\rm{Q,41}}^{-\frac{1}{2}}\frac{P_{\rm{1}}}{M_{\rm{6}}}\eta_{\rm{-1}}^{\frac{1}{2}},$ (13)
where $L_{\rm{Q,41}}=L_{\rm{Q}}/\left(10^{41}\ \rm{erg/s}\right)$ and $L_{\rm{p,42}}=L_{\rm{p}}/\left(10^{42}\right.$ $\left.\rm{erg/s}\right)$. In this way, we obtain two constraints on $r_{*}$, both of which must be satisfied, and each contains one fudge factor only.
3.7 Radiation Temperature
@@
When the star-disk collision occurs, the material shocked by the collision will be radiation-dominated and optically thick. Firstly, the radiation energy density is expressed as $u_{\gamma}=L_{\rm{p}}/[4\pi R_{\rm{diff}}^{2}v_{\rm{ej}}]$ (Linial & Metzger, 2023). Secondly, assuming that the radiation of QPEs is dominated by blackbody radiation, $u_{\gamma}=aT_{\rm{p}}^{4}$, where $a$ is the radiation constant, then $T_{\rm{p}}$ can be expressed in terms of various observables as (also see Linial & Metzger, 2023)
$kT_{\rm{p}}$ $=k\left(\frac{u_{\gamma}}{a}\right)^{\frac{1}{4}}$
$=9\ L_{\rm{p,42}}^{\frac{1}{4}}\left(\frac{t_{\rm{e}}}{1\ \rm{hr}}\right)^{-\frac{1}{2}}P_{1}^{\frac{1}{4}}M_{\rm{6}}^{-\frac{1}{4}}\ \rm{eV},$ (14)
@@
Note that Linial & Metzger (2023) argue that if the radiation is not in thermal equilibrium with the matter at $R_{\rm{diff}}$, then $kT_{\rm{p}}$ might be higher.
3.8 Constraint from Stellar Tidal Disruption
Now we consider another constraint on $r_{*}$. The stellar orbit must be sufficiently wide; otherwise the star would be tidally disrupted. A critical orbital radius is the so-called tidal disruption radius (e.g., Rees, 1988)
$R_{\rm{t}}=R_{*}\left(\frac{M_{\rm{h}}}{M_{*}}\right)^{\frac{1}{3}}=47r_{*}M_{\rm{6}}^{-\frac{2}{3}}m_{*}^{-\frac{1}{3}}\ R_{\rm{g}}.$ (15)
The above constraint requires that $R>nR_{\rm{t}}$, where $n=1$ corresponds to the full disruption and $n=2$ corresponds to the starting point of mass transfer; In other words, $21<n<$ 2 corresponds to the partial disruption of the star.
Plugging Eq. (1) in, this can be rearranged to put a constraint on the stellar radius: $r_{*}<2.4\ n^{-1}P_{\rm{1}}^{\frac{2}{3}}m_{*}^{\frac{1}{3}}$. Adopting the mass-radius relationship for high-mass zero-age-main-sequence stars $r_{*}=m_{*}^{0.6}$ (Kippenhahn et al., 2013), one can rewrite it as
$r_{\rm{*}}<6.9\ n^{-\frac{9}{4}}P_{1}^{\frac{3}{2}}.$ (16)
Figure 1: The constraints on the stellar radius for GSN 069 (left panel) and RX J1301 (right panel). The red and blue lines stand for the constraints based on Eq. (12) and Eq. (13), respectively. The solid and dashed lines are for different adopted values of $0.1$ and $0.01$ for the model parameter ($\alpha$ or $\eta$), respectively. The horizontal lines with downward arrows represent the full tidal disruption limit (green, Eq. 16, $n=1$) and the partial tidal disruption limit (black, Eq. 16, $n=2$), respectively.
@@ 4 Application to QPE Sample
In this section, for each source, we plot the $r_{*}$-constraints as functions of $M_{\rm{h}}$. The numerical range of $M_{\rm{h}}$ of an individual source corresponds to the estimate quoted in Sect. 2 and Table. 1.
4.1 GSN 069
@@
We start with the well-known QPE source GSN 069. In Figure 1, we plot the $r_{*}$-constraints, Eq. (12) and Eq. (13) as the red and blue lines, respectively. The solid and dashed lines are for two adopted representative values of the unknown parameters ($\alpha$ and $\eta$), respectively. The light-purple shadowed region represents the allowed parameter space that could satisfy both Eq. (12) and Eq. (13). The full tidal disruption limit and the partial tidal disruption limit are represented by the green and black horizontal lines, respectively. The region below the green line means that stars with such small $r_{*}$ can avoid being tidally disrupted. Similarly, stars with even smaller $r_{*}$ (below the black line) will not even be tidally stripped.
@@
As the left panel of Figure 1 shows, GSN 069 does have a light-purple shaded region of $r_{*}\lesssim 10$, especially for $M_{\rm{h}}\lesssim 10^{6}\ M_{\rm{\odot}}$, that satisfies both constraints. However, most of that region is above the solid green line, meaning that they would have already been tidally disrupted.
@@
Only a small shaded region with small $r_{*}$ (i.e., below the green solid line) can avoid the full disruption, but they are still well above the black solid line, meaning that they would suffer a significant partial disruption and that disruption would repeat on every orbit. In the latter case, the question arises that whether the stripped mass and its subsequent accretion toward the BH will be the dominant energy source for the observed eruption. Indeed, a few groups have argued for the repeated strippings or mass transfer from a stellar companion as the origin of QPEs (King, 2020, 2022, 2023; Zhao et al., 2022; Chen et al., 2023; Wu et al., 2024; Krolik & Linial, 2022; Lu & Quataert, 2023).
In any case, these results demonstrate that the simplest version of the star-disk collision model did not pass our test for GSN 069.
4.2 Other QPE Sources

### v3

@@ III Constraints for the Model
$R=\left[GM_{\rm{h}}\left(\frac{P_{\rm{orb}}}{2\pi}\right)^{2}\right]^{\frac{1}{3}}=110P_{\rm{1}}^{\frac{2}{3}}M_{\rm{6}}^{-\frac{2}{3}}\ R_{\rm{g}},$ (1)
where $R_{\rm{g}}=\frac{GM_{\rm{h}}}{c^{2}}$ is the gravitational radius of the BH and $c$ is the speed of light in vacuum. In order to simplify the derivation, we define the following notation: $r_{\star}=R_{\star}/\rm{R}_{\rm{\odot}}$, $M_{\rm{6}}=M_{\rm{h}}/10^{6}\ \rm{M}_{\rm{\odot}}$, $P_{\rm{1}}=P_{\rm{orb}}/10\ \rm{hr}$ and $\alpha_{-1}=\alpha/0.1$, where $\rm{R}_{\rm{\odot}}$ and $\rm{M}_{\rm{\odot}}$ are the solar radius and mass, $\alpha$ is the dimensionless disk viscosity parameter and $R_{\star}$ is the radius of the star.
III.1 The Disk
@@
For the inner region of the disk where it is radiation-pressure and radiative-cooling dominated, the scale height of the disk is given by (also see Frank et al., 2002; Linial and Metzger, 2023),
@@
where $v_{\rm{k}}=\sqrt{\frac{GM_{\rm{h}}}{R}}$ is the Kepler velocity at radius $R$, $\kappa$ is the opacity, $\dot{M}$ is the mass accretion rate at the collision point. Hereafter, we consider that the opacity is dominated by Thompson scattering, thus $\kappa=0.34\ \rm{cm}^{2}\rm{g}^{-1}$. The last step of the above has used $R\gg 2R_{\rm{g}}$. We use the lowercase letter $\dot{m}\equiv\dot{M}/\left(L_{\rm{Edd}}/c^{2}\right)$ to denote the dimensionless accretion rate, where $L_{\rm{Edd}}$ is the Eddington luminosity. Then, the luminosity created by accretion is $L=\eta\dot{m}L_{\rm{Edd}}$, where $\eta$ is the radiation efficiency.
Based on the $\alpha$-disk model (Shakura and Sunyaev, 1973), the surface density is given as
@@
where $\nu\approx\alpha c_{\rm{s}}h\approx\alpha\sqrt{GMR}(\frac{h}{R})^{2}$ is the viscosity of the disk.
@@
$M_{\rm{sh}}$ $=2\pi R_{\star}^{2}\left(\sqrt{2}\Sigma_{\rm{d}}\right)$
$=4.3\times 10^{-7}r_{\star}^{2}\frac{P_{\rm{1}}}{\alpha_{-1}\dot{m}M_{6}}\ \rm{M}_{\rm{\odot}}.$ (4)
The factor of $\sqrt{2}$ comes from the consideration that the velocity of the disk material is perpendicular to the stellar velocity and both are of the order of $v_{\rm{k}}$, so the effective length of the stellar trajectory is $2\sqrt{2}h$.
III.2 Hydrodynamics in Collision
@@ III.3 Duration of Emission
$\Sigma=\frac{M_{\rm{sh}}}{4\pi R_{\rm{diff}}^{2}}$ (7)
is the surface density of the fireball, $R_{\rm{diff}}=v_{\rm{ej}}t_{\rm{e}}$ is the radius of the fireball at its peak and $t_{\rm{e}}$ is the duration of the eruption. Therefore, one has (also see Linial and Metzger, 2023)
$t_{\rm{e}}\approx\left(\frac{\kappa M_{\rm{sh}}}{4\pi cv_{\rm{ej}}}\right)^{\frac{1}{2}}\approx 0.18\frac{r_{\star}}{\alpha_{-1}^{\frac{1}{2}}\dot{m}^{\frac{1}{2}}}\left(\frac{P_{\rm{1}}}{M_{\rm{6}}}\right)^{\frac{2}{3}}\ \rm{hr}.$ (8)
III.4 Luminosity
Next, we consider the energy evolution of the shocked matter. The following Eq. (5), the total internal energy of the immediately post-shock matter, which is dominated by radiation, is $E_{\rm{i,0}}=M_{\rm{sh}}v_{\rm{k}}^{2}$. Note that it is on the same order of magnitude as the total kinetic energy of the shocked matter $E_{\rm{k}}=\frac{1}{2}M_{\rm{sh}}v_{\rm{k}}^{2}$ (Linial and Metzger, 2023).
@@ III.4 Luminosity
$E_{i}\left(R_{\rm{diff}}\right)$ $=E_{\rm{i,0}}\times\left(\frac{2\pi R_{\star}^{2}\sqrt{2}h/7}{\frac{4\pi}{3}R_{\rm{diff}}^{3}}\right)^{\frac{1}{3}}$
$=3.9\times 10^{44}r_{\star}^{\frac{5}{3}}\dot{m}^{-\frac{1}{6}}\alpha_{\rm{-1}}^{-\frac{1}{2}}M_{\rm{6}}^{\frac{1}{3}}\ \rm{erg/s}.$ (9)
@@
$L_{\rm{p}}=\frac{E_{i}\left(R_{\rm{diff}}\right)}{t_{\rm{e}}}=6.1\times 10^{41}M_{\rm{6}}\left(\frac{r_{\star}}{P_{\rm{1}}}\right)^{\frac{2}{3}}\dot{m}^{\frac{1}{3}}\ \rm{erg/s}.$ (10)
@@
III.5 The Quiescent-level Accretion
@@
III.6 Infer $r_{\star}$ from Observables
Eqs. (8), (10) and (11) give the three observables $(t_{\rm{PQE}}$, $L_{\rm{p}}$ and $L_{\rm{Q}})$ from the shock-cooling emission model induced by star-disk collisions. Here, we attempt to investigate what $r_{\star}$ is allowed or required by the model to satisfy these observables.
Notice that $\alpha$ and $\eta$ are the two fudge “factors” that describe the working mechanism of the accretion disk, and $\dot{m}$ is one of the disk properties; all others are observable, or measurable in principle. One can get rid of $\dot{m}$ and express $r_{\star}$ in terms of observables in two ways. First, one can infer $\dot{m}$ from Eq. (10), and plug it into Eq. (8) to get
$r_{\star}=4L_{\rm{p,42}}^{\frac{3}{4}}P_{\rm{1}}^{\frac{1}{6}}M_{\rm{6}}^{-\frac{5}{12}}\left(\frac{t_{\rm{e}}}{1\ \rm{hr}}\right)^{\frac{1}{2}}{\alpha_{-1}}^{\frac{1}{4}}.$ (12)
@@ III.6 Infer $r_{\star}$ from Observables
$r_{\star}=32\ L_{\rm{p,42}}^{\frac{3}{2}}L_{\rm{Q,41}}^{-\frac{1}{2}}\frac{P_{\rm{1}}}{M_{\rm{6}}}\eta_{\rm{-1}}^{\frac{1}{2}},$ (13)
where $L_{\rm{Q,41}}=L_{\rm{Q}}/\left(10^{41}\ \rm{erg/s}\right)$ and $L_{\rm{p,42}}=L_{\rm{p}}/\left(10^{42}\right$. $\left.\rm{erg/s}\right)$. In this way, we obtain two constraints on $r_{\star}$, both of which must be satisfied, and each contains one fudge factor only.
III.7 Radiation Temperature
@@
When the star-disk collisions occur, the material shocked by the collision will be radiation-dominated and optically thick. Firstly, the radiation energy density is expressed as $u_{\gamma}=L_{\rm{p}}/[4\pi R_{\rm{diff}}^{2}v_{\rm{ej}}]$ (Linial and Metzger, 2023). Secondly, assuming that the radiation of QPEs is dominated by blackbody radiation, $u_{\gamma}=aT_{\rm{p}}^{4}$, where $a$ is the radiation constant, then $T_{\rm{p}}$ can be expressed in terms of various observables as (also see Linial and Metzger, 2023)
$kT_{\rm{p}}=9\ L_{\rm{p,42}}^{\frac{1}{4}}\left(\frac{t_{\rm{e}}}{1\ \rm{hr}}\right)^{-\frac{1}{2}}P_{1}^{\frac{1}{4}}M_{\rm{6}}^{-\frac{1}{4}}\ \rm{eV},$ (14)
@@
Note that Linial and Metzger (2023) argue that if the radiation is not in thermal equilibrium with the matter at $R_{\rm{diff}}$, then $kT_{\rm{p}}$ might be higher.
III.8 Constraint from Stellar Tidal Disruption
Now we consider another constraint on $r_{\star}$. The stellar orbit must be sufficiently wide; otherwise the star would be tidally disrupted. A critical orbital radius is the so-called tidal disruption radius (e.g., Rees, 1988)
$R_{\rm{t}}=R_{\star}\left(\frac{M_{\rm{h}}}{M_{*}}\right)^{\frac{1}{3}}=47r_{\star}M_{\rm{6}}^{-\frac{2}{3}}m_{*}^{-\frac{1}{3}}\ R_{\rm{g}}.$ (15)
The above constraint requires that $R>nR_{\rm{t}}$, where $n=1$ corresponds to the full disruption and $n=2$ corresponds to the starting point of mass transfer; In other words, $1<n<2$ corresponds to the partial disruption of the star.
Plugging Eq. (1) in, this can be rearranged to put a constraint on the stellar radius: $r_{\star}<2.4\ n^{-1}P_{\rm{1}}^{\frac{2}{3}}m_{*}^{\frac{1}{3}}$. For a normal star, one can adopt the mass-radius relationship for high-mass zero-age-main-sequence stars : $r_{\star}=m_{*}^{0.6}$ (Kippenhahn et al., 2013). However, as was mentioned in Linial and Metzger (2023); Linial and Sari (2023); Linial and Quataert (2024) and was shown in simulation by Yao et al. (2025), the star may get heated and inflated during collisions. The increased $r_{\star}$ would cause $R_{\rm{t}}$ in Eq. (15) to be larger, making it easier to be disrupted. Then the condition of $R>nR_{\rm{t}}$ would exclude a boarder parameter space.
To include this effect, we introduce a bloating factor $b$ into the mass-radius relations for an inflated star : $r_{\star}=bm_{*}^{0.6}$. Then, one can rewrite the condition for avoiding the disruption as
$r_{\rm{*}}<6.9\ n^{-\frac{9}{4}}P_{1}^{\frac{3}{2}}b^{-\frac{5}{4}}.$ (16)
Figure 1: The constraints on the stellar radius for GSN 069 (left panel) and RX J1301 (right panel). The red and blue lines stand for the constraints based on Eq. (12) and Eq. (13), respectively. The solid and dashed lines are for different adopted values of $0.1$ and $0.01$ for the model parameter ($\alpha$ or $\eta$), respectively. The horizontal lines with downward arrows represent the full tidal disruption limit (green, Eq. 16, $n=1$ and $b=1$) and the partial tidal disruption limit (black, Eq. 16, $n=2$ and $b=1$), respectively.
@@ IV Application to QPE Sample
In this section, for each source, we plot the $r_{\star}$-constraints as functions of $M_{\rm{h}}$. The numerical range of $M_{\rm{h}}$ of an individual source corresponds to the estimate quoted in Sect. II and Table. 1.
IV.1 GSN 069
@@
We start with the well-known QPE source GSN 069. In Figure 1, we plot the $r_{\star}$-constraints, Eq. (12) and Eq. (13) as the red and blue lines, respectively. The solid and dashed lines are for two adopted representative values of the unknown parameters ($\alpha$ and $\eta$), respectively. The light-purple shadowed region represents the allowed parameter space that could satisfy both Eq. (12) and Eq. (13). The full tidal disruption limit and the partial tidal disruption limit are represented by the green and black horizontal lines, respectively. The region below the green line means that stars with such small $r_{\star}$ can avoid being tidally disrupted. Similarly, stars with even smaller $r_{\star}$ (below the black line) will not even be tidally stripped.
@@
As the left panel of Figure 1 shows, GSN 069 does have a light-purple shaded region of $r_{\star}\lesssim 10$, especially for $M_{\rm{h}}\lesssim 10^{6}\ M_{\rm{\odot}}$, that satisfies both constraints. However, most of that region is above the solid green line, meaning that they would have already been tidally disrupted.
@@
Only a small shaded region with small $r_{\star}$ (i.e., below the green solid line) can avoid the full disruption, but they are still well above the black solid line, meaning that they would suffer a significant partial disruption and that disruption would repeat on every orbit. In the latter case, the question arises that whether the stripped mass and its subsequent accretion toward the BH will be the dominant energy source for the observed eruption. Indeed, a few groups have argued for the repeated strippings or mass transfer from a stellar companion as the origin of QPEs (King, 2020, 2022, 2023; Zhao et al., 2022; Chen et al., 2023; Wu et al., 2024; Krolik and Linial, 2022; Lu and Quataert, 2023). These results demonstrate that the simplest version of the shock-cooling emission model from star-disk collisions did not pass our test for GSN 069.
IV.2 Other QPE Sources